I am having trouble with probability. If the odds are 2:1 then the total chances is 3 and chances for is 2 and chances against is 1. is that correct? That makes the odds ratio .67/.33 or 2. (chances for / chances against).
if the probability is .18 then .82 is chances against or the odds are .22:1. Total chances is 100
is that correct? My biggest problem is determining odds against and odds in favor. Somehow I am getting this all mixed up.
Thanks
Randy

Answer by Edwin McCravy(20081)




                     Number of successes
Probability = —————————————————————-------------------
               Number of success + Number of failures 


Odds in favor = Number of successes : Number of failures

[Read the colon : as the word "to"]

Odds against = Number of failures : Number of successes

Ofentimes you will see Odds expressed so that it is "something to one"

To express it "to one" form:

                 Number of successes     
Odds in favor = ————————————————————— : 1
                 Number of failures 


                Number of failures     
Odds against = ———————————————------ : 1
                Number of successes 


So the abreviations for the above formulas:
      S
P = —————
     S+F 

Odds in favor = S:F

Odds against = F:S

if the probability is .18

then .18 = 18/100 = 9/50 so S = 9 and S+F = 50

So F = 50-S = 50-9 = 41

So Odds in favor = S:F or 9:41

Odds against = 41:9

Or you can divide 41 and 9 by 9 and get

Odds against = 4.6 to 1.

Edwin